OpenStudy (anonymous):

Write the equation of a line that passes through the point (-1, 2) with a slope of - 5 2 in slope-intercept.

OpenStudy (igreen):

Your equation will be \(y = -\dfrac{5}{2}x - \dfrac{1}{2}\) because it crosses through (-1, 2) and the slope of the equation(m) is \(-\dfrac{5}{2}\).

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @ryanmar